OpenStudy (anonymous):

for the love of god ive been up all night someone please just tell me the answers to each of these -2<2x-4<4 2|3x-5|-8=8 6|6-4x|=8x +4 8|x+ 30ver4|<2

OpenStudy (anonymous):

the last one is supposed to say 3 over 4

OpenStudy (anonymous):

\[6\left| 6-4x \right|=8x+4\] \[\left| 6-4x \right|=\frac{ 8x+4 }{ 6 }\] \[6-4x=\pm \frac{ 8x+4 }{ 6 }\]

OpenStudy (anonymous):

whats the answer haha

OpenStudy (anonymous):

\[either 6-4x=\frac{ 8x+4 }{6 }\] 6(6-4x)=8x+4 36-24x=8x+4 36-4=8x+24x 32x=32,x=32/32=1

OpenStudy (anonymous):

ok that was correct what are the other answers

OpenStudy (anonymous):

\[or6-4x=-\frac{ 8x+4 }{6 }\] 6(6-4x)=-(8x+4) 36-24x+8x+4=0 36+4=24x-8x 16x=40 x=40/16=5/2

OpenStudy (anonymous):

answer my other questions not just that one

OpenStudy (anonymous):

i solved a bit difficult ,others ar easy. you start i will guide you.

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

let us start with first. write the statement first.

OpenStudy (anonymous):

then add 4 in all the three inequalities.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

not like this -2+4<2x-4+4<4+4 2<2x<8 now divide each inequality by 2

OpenStudy (anonymous):

1<x<4

OpenStudy (anonymous):

how you got 6 ?

OpenStudy (anonymous):

now write the second statement and show me.

OpenStudy (anonymous):

\[2\left| 3x-5 \right|-8=8\] \[2\left| 3x-5 \right|=8+8\] \[\left| 3x-5 \right|=\frac{ 16 }{2 }=8\] \[3x-5=\pm8\]

OpenStudy (anonymous):

either 3x-5=8 or 3x-5=-8 now you can solve these equations for x.

OpenStudy (anonymous):

are you solving ?

OpenStudy (anonymous):

\[8\left| x+\frac{ 3 }{4 } \right|<2\] \[\left| x+\frac{ 3 }{4 } \right|<\frac{ 2 }{8}\] \[\left| x+\frac{ 3 }{ 4 } \right|<\frac{ 1 }{4 }\] \[-\frac{ 1 }{ 4 }<x+\frac{ 3 }{4 }<\frac{ 1 }{4 }\] multiply by 4 -1<4x+3<1 subtract 3 -1-3<4x+3-3<1-3 -4<4x<-2 divide by 4 \[-1 <x<\frac{ -2 }{ 4 },or-1<x<-\frac{ 1 }{2 }\]